∫ (x)= ∫ (sinxy) dx+ ∫ (cotxy) dx


1
Expert's answer
2020-04-10T17:53:14-0400

"\int sin(xy)dx+\int cot(xy)dx=1/y\int sin(xy)d(xy)+1/y\int cot(xy)d(xy)=\\-cos(xy)/y+ln(sin(xy))/y+C=(ln(sin(xy))-cos(xy))/y+C"